# MetaSel: A Test Selection Approach for Fine-tuned DNN Models

### v2

@@ IV-A Determining MetaSel’s feature set
To identify an effective set of input features $X_{m}$ for training a model accurately predicting misclassified inputs, we conducted a preliminary study exploring various feature sets. We considered features extracted from the target training inputs (images in our experiments), the final class probabilities of the source ($M_{S}$) and target ($M_{T}$) models, and their output labels, as well as a measure of how closely an input conforms to the data distribution on which the source and target DNN models were trained. For the latter, we employed ODIN (Out-of-DIstribution detector for Neural networks) [18], a highly effective yet efficient approach that achieves robust Out-of-Distribution (OOD) detection without requiring any architectural modifications or additional training. It is crucial for OOD detection in MetaSel to be lightweight to maintain its efficiency while ensuring its effectiveness. Based on the results of our preliminary study and the reasons discussed below, we considered the following features, as illustrated in Figure 2:
  • The output of the logit layer of $M_{S}$ ($L_{S}$),
  • The output of the logit layer of $M_{T}$ ($L_{T}$),
  • The difference between $L_{S}$ and $L_{T}$,
@@
At a high level, the assumptions underlying these features include (1) The more uncertain the predictions, the more likely they are to be incorrect, (2) If the two models’ predictions are inconsistent, then the corresponding inputs lie in the input subspace where the decision boundary was changed as a result of fine-tuning and are more likely to be misclassified, (3) OOD inputs are more likely to be misclassified. Since $M_{T}$ is a version of $M_{S}$ specifically tailored to the target context, the two models share a foundational understanding, through the knowledge transferred during fine-tuning. Thus, information from both models is potentially relevant.
@@
We leverage logits since they allow for more fine-grained comparisons between the two DNN models and constitute more precise measurements of inconsistency or disagreement between predictions than the final class probabilities or output labels alone. Compared to the output label for models $M_{S}$ and $M_{T}$, using the logit layers offers several advantages. Labels are determined by selecting the class with the highest probability from the probability distribution produced by the last layer’s activation function (softmax layer). In contrast, logits represent the raw scores produced by the layer preceding the softmax layer. Therefore, unlike the output label which only indicates the predicted class, logits reflect the model’s raw scoring for each output class, providing a more complete view of the model’s internal decision-making. Consequently, logits provide a richer representation of the pre-trained and fine-tuned models’ behavior and a finer indication of their agreement or disagreement. For instance, consider the case where both the pre-trained and fine-tuned models produce the same output class for a given input, while one model produced logits indicating high confidence in its prediction and the other produced logits showing very close values for two or more classes, reflecting uncertainty. In this scenario, the output labels alone suggest agreement between the two models whereas their logits reveal discrepancies in their prediction confidence. Consequently, we rely on logits as they provide a richer representation of the model’s behavior compared to output labels.
@@
The output of the logit layer in a DNN represents raw, unnormalized scores. These scores are then passed through a softmax activation function, which converts them into class probabilities. Class probabilities are normalized values of logits, restricted within the range of [0, 1], and collectively add up to one. However, this normalization procedure can potentially obscure the information originally contained in the logits. Indeed, Aigrain et al. [19] further demonstrated that logits have greater discriminative power than softmax values, making the former more effective in detecting misclassifications and adversarial examples in DNNs. Additionally, the normalization procedure amplifies the differences between logits, making class probabilities vulnerable to outliers present in the logits. In other words, outliers (i.e., extremely high or low logits), can disproportionately affect the output probabilities. In extreme cases, a single outlier can dominate the probability distribution, with its associated class probability approaching 1, while the probabilities for all other classes converge to 0.
@@
However, to ensure that we use all the relevant information available, we also include in our feature set a boolean value indicating whether the output labels of $M_{S}$ and $M_{T}$ are identical or differ. Furthermore, for the reasons presented above, we also feed MetaSel the difference between the logits of two models for each output class, as this may enhance its discriminative power. The difference between $L_{S}$ and $L_{T}$ indicates the magnitude of discrepancies between the prediction of the pre-trained and fine-tuned models for each output class, indicating whether the given input is near the input subspace where the decision boundary has changed.
@@
Notably, we also observed that incorporating representations of target input images derived from a pre-trained VGG-16 [20] model in MetaSel’s feature set did not improve its performance. This may be explained by the fact that information about a given input image is already encoded in the outputs of the logit layer [21].
As part of our preliminary study, we conducted an ablation study to validate the usefulness of the selected features presented above. Results of this study, reported in our replication package [22], confirm the contribution of each individual feature to the overall performance of MetaSel.
@@ IV-B Constructing MetaSel’s training set
To train MetaSel, we require a set of inputs from the target distribution ($D_{T}$) whose corresponding MetaSel’s output are known. To be able to determine MetaSel’s output, we need to know the following information for each input ($t\in X_{T}$):
@@
To address this, we employed once again ODIN scores [18], to filter out source test inputs that do not align with the target data distribution, $D_{T}$. Source test inputs with ODIN scores lower than a specific threshold are considered OOD and filtered out. To determine an optimized threshold for each $M_{T}$, we rely on the common procedure introduced by Liang et al. [18] that involves using a separate validation set. Following this procedure, we select a threshold that minimizes the False Positive Rate (FPR) at a fixed True Positive Rate (TPR) of 95%. This threshold ensures that the model retains a high level of accuracy in distinguishing ID inputs from OOD inputs. By applying this filtering mechanism, as illustrated in Figure 2, we ensured that only those source test inputs that satisfy both criteria were included in the MetaSel training set.
@@ V-B Subjects
We conducted our experiments using a collection of well-known image input sets, including MNIST [23], Fashion-MNIST [23], Cifar-10 [24], and Cifar-100 [24], that have been utilized in numerous empirical studies on DNN testing [25, 2, 26, 17, 27]. MNIST is a set of images of handwritten digits and Fashion-MNIST (FMNIST) contains images of clothes that are associated with fashion and clothing items. Each one represents a 10-class classification problem and consists of 70,000 grayscale images of a standardized 28x28 size. The training set of each input set includes 60,000 images, while each test set includes 10,000 images. Cifar-10 is another widely-used input set that includes a collection of images from 10 different classes (e.g., cats, dogs, airplanes, cars). Furthermore, we selected a more complex input set, Cifar-100, containing 100 classes with 60,000 images total, known for its realistic and feature-rich images. Both Cifar-10 and Cifar-100 are widely used benchmarks in computer vision and contain 32x32 cropped colored images. Each input set is divided into 50,000 images for training and 10,000 images for testing.
@@
In addition to corrupted test sets, we require corrupted versions of our source training sets, i.e., target training set, to perform the fine-tuning process and obtain fine-tuned models. However, Cifar-10-C, Cifar-100-C, MNIST-C, and FMNIST-C datasets only provide corrupted test sets. Therefore, we utilized the original source code and settings employed in the creation of these corrupted test sets to generate the corrupted versions of our source training sets. However, in practical scenarios, the size of the target training set is often limited and significantly smaller than the source training set due to labeling costs. Therefore, we sampled a subset of these corrupted training sets to serve as the target training set, as detailed in the following section. We refer to each fine-tuned DNN model ($M_{T}$) and its corresponding training and test input sets ($Train_{T}$ and $Test_{T}$, respectively) as subjects in our experiments. Consequently, we conducted a comprehensive study across 68 subjects, encompassing four source input sets, seven corruption types, and three levels of corruption severity.
@@ V-D Evaluation Measure
[…]
where $Mis_{total}$ represents the total number of failures (i.e., mispredicted inputs) in the entire test set, $Mis_{T}$ is the number of failures detected by test subset $T$, and $b$ denotes the selection budget, i.e., size of $T$. Note that the denominator in the TRC formula, as defined in Equation (5), adjusts for budget sizes smaller than $Mis_{total}$. This ensures that TRC accurately evaluates how effective test selection is compared to the ideal performance within the given budget. If the denominator were fixed at $Mis_{total}$, regardless of the budget, achieving a TRC value of 1 would be impossible for any subset selected with $b<Mis_{total}$. For example, with $Mis_{total}$=20 and a selection budget of $b$=10, no selected subset could reach a TRC value of 1 due to the inherent budget constraint. The formulation in Equation (5) overcomes this limitation by adapting to the budget, thereby providing a meaningful and general measure of effectiveness across all budgets and regardless of the number of mispredicted inputs.
@@
Figure 3: TRC values for MetaSel and baseline approaches across diverse selection budgets on all subjects with a medium severity level of distribution shift (severity level = 3)
@@
Finally, our primary focus in this context is on evaluating an approach’s capability to identify the maximum number of misclassifications within a specified selection budget. The TRC metric effectively captures this objective. Therefore, we employ TRC in our experiments to assess and compare the performance of MetaSel against the baseline methods.
@@ V-E Execution Environment
We conducted our experiments using a cloud computing environment provided by the Digital Research Alliance of Canada [36], utilizing the Cedar cluster with 1352 various Nvidia GPUs (P100 Pascal and V100 Volta) with memory ranging from 12 to 32 GB. Conducting our extensive experiment, which involved deploying MetaSel and all baselines across all subjects and distribution shift severity levels, required approximately 39 days to complete.
V-F Data Availability
The replication package for our studies will be shared upon the paper’s acceptance [22].
@@ VI-A RQ1: MetaSel’s Performance improvement
Using a simpler improvement percentage formula111$\left(\frac{TRC_{MetaSel}(b)-TRC_{bestBaseline}(b)}{TRC_{bestBaseline}(b)}\right)\times 100$, solely based on the TRC value achieved by baselines, might overstate smaller gains. Consequently, the improvement percentage calculated based on Equation (7) ensures a more accurate analysis of MetaSel’s effectiveness relative to baselines. Furthermore, compared to calculating absolute improvements, this formula normalizes improvements and allows for a meaningful comparison across budgets and subjects.
@@
Both DATIS and NNS demonstrate significantly lower TRC medians compared to MetaSel. They also show greater variability in TRC values, introducing the risk of ineffective test selection for some subjects. These findings confirm MetaSel’s superiority and its ability to maintain consistently higher TRC under tight selection budgets, making it the most effective selection approach among all evaluated methods. Moreover, the results highlight the risk of relying on alternative baselines, even the most frequent second-best approaches, like Meta-model, DATIS, or NNS, as they may lead to significantly poor performance.
@@
Gini Margin Vanilla DSA MDSA LSA TestRank Meta-model DATIS NNS
LeNet5 MNIST 10.45 1.16 0.95 0.74 165.07 1.08 5.40 79.83 9.23 8.98 5.04
LeNet5 FMNIST 10.64 1.15 0.99 0.97 172.69 1.09 11.03 97.29 11.69 10.11 4.85
ResNet20 Cifar-10 51.02 2.14 2.02 2.01 167.40 2.98 9.22 92.46 24.81 9.74 5.55
ResNet152 Cifar-100 277.95 27.14 25.65 25.58 228.40 48.32 63.32 118.07 382.05 91.41 83.25
@@ VI-B RQ2: MetaSel’s efficiency
While MetaSel, similar to many baseline approaches, exhibits an increase in execution time as the complexity of the DNN model and the size of the test set grow, it is not the most time-consuming approach. When applied to subjects with relatively simple architectures, such as LeNet-5 or ResNet-20, and input sets with only 10 output classes (i.e., MNIST, FMNIST, or Cifar-10), MetaSel demonstrates efficiency outperforming baselines such as TestRank and DSA in terms of execution time. The primary factors driving MetaSel’s execution time are the depth and complexity of the DNN model as well as the number of output classes. This is primarily due to the requirement for ODIN score calculations for test inputs based on both the pre-trained and fine-tuned models, which requires significantly more computations for forward and backward passes to calculate the ODIN scores on deeper networks such as ResNet152. Additionally, the size of the logit layer, which corresponds to the number of output classes, further increases computational requirements for input sets like Cifar-100. Despite these challenges, MetaSel maintains acceptable efficiency even when applied to deeper networks like ResNet152 and input sets with a larger number of output classes, such as Cifar-100 with 100 classes. Indeed, MetaSel requires approximately 278 seconds to rank the entire test set, a remarkably lower execution time compared to Meta-model, the most viable alternative for practitioners based on effectiveness, which takes around 382 seconds. The increased execution time, combined with the longer initial training phase required by Meta-model, highlights the considerably more time-intensive nature of this method compared to MetaSel.
@@
Answer to RQ2: MetaSel’s efficiency, though lower than many baselines, remains acceptable from a practical standpoint, even for subjects with large input sets featuring numerous output classes, and deep, complex DNN architectures. Its efficiency is even better than the second-best baseline in terms of effectiveness, namely Meta-model. MetaSel is thus a cost-effective approach.

### v4

@@ IV-A Determining MetaSel’s feature set
To identify an effective set of input features $X_{m}$ for training a model accurately predicting misclassified inputs, we conducted a preliminary study exploring various feature sets. We considered features extracted from the target training inputs (images in our experiments), the final class probabilities of the source ($M_{S}$) and target ($M_{T}$) models, and their output labels, as well as a measure of how closely an input conforms to the data distribution on which the source and target DNN models were trained. For the latter, we employed ODIN (Out-of-DIstribution detector for Neural networks) [23], a highly effective yet efficient approach that achieves robust Out-of-Distribution (OOD) detection without requiring any architectural modifications or additional training.
While MetaSel can be configured to deploy any OOD detection approach, it is crucial to use lightweight solutions to maintain MetaSel’s efficiency while ensuring its effectiveness. For instance, Likelihood Regret (LR) OOD [24] has demonstrated higher accuracy than ODIN across a broad range of input sets [24] and could potentially enhance MetaSel’s effectiveness. However, LR OOD entails significantly higher computational cost, which would make MetaSel less efficient, particularly when applied to larger input sets or complex DNNs. Based on the results of our preliminary study and the reasons discussed below, we considered the following features, as illustrated in Figure 2:
  • The output of the logit layer of $M_{S}$ (vector $L_{S}$),
  • The output of the logit layer of $M_{T}$ (vector $L_{T}$),
  • A vector representing the element-wise absolute difference between $L_{S}$ and $L_{T}$,
@@
At a high level, the assumptions underlying these features include (1) The more uncertain the predictions, the more likely they are to be incorrect, (2) If the two models’ predictions are inconsistent, then the corresponding inputs lie in the input subspace where the decision boundary was changed as a result of fine-tuning and are more likely to be misclassified, (3) OOD inputs are more likely to be misclassified. Since $M_{T}$ is a version of $M_{S}$ specifically tailored to the target context, the two models share a foundational understanding through the knowledge transferred during fine-tuning. Thus, information from both models is potentially relevant.
@@
We leverage logits since they allow for more fine-grained comparisons between the two DNN models and constitute more precise measurements of inconsistency or disagreement between predictions than the final class probabilities or output labels alone. Compared to the output label for models $M_{S}$ and $M_{T}$, using the logit layers offers several advantages. Labels are determined by selecting the class with the highest probability from the probability distribution produced by the last layer’s activation function (softmax layer). In contrast, logits represent the raw scores produced by the layer preceding the softmax layer. Therefore, unlike the output label, which only indicates the predicted class, logits reflect the model’s raw scoring for each output class, providing a more complete view of the model’s internal decision-making. Consequently, logits provide a richer representation of the pre-trained and fine-tuned models’ behavior and a finer indication of their agreement or disagreement. For instance, consider the case where both the pre-trained and fine-tuned models produce the same output class for a given input, while one model produces logits indicating high confidence in its prediction and the other produces logits showing very close values for two or more classes, reflecting uncertainty. In this scenario, the output labels alone suggest agreement between the two models, whereas their logits reveal discrepancies in their prediction confidence. Consequently, we rely on logits as they provide a richer representation of the model’s behavior compared to output labels.
@@
The output of the logit layer in a DNN represents raw, unnormalized scores. These scores are then passed through a softmax activation function, which converts them into class probabilities. Class probabilities are normalized values of logits, restricted within the range of [0, 1], and collectively add up to one. However, this normalization procedure can potentially obscure the information originally contained in the logits. Indeed, Aigrain et al. [25] further demonstrated that logits have greater discriminative power than softmax values, making the former more effective in detecting misclassifications and adversarial examples in DNNs. Additionally, the normalization procedure amplifies the differences between logits, making class probabilities vulnerable to outliers present in the logits. In other words, outliers (i.e., extremely high or low logits) can disproportionately affect the output probabilities. In extreme cases, a single outlier can dominate the probability distribution, with its associated class probability approaching 1, while the probabilities for all other classes converge to 0.
@@
However, to ensure that we use all the relevant information available, we also include in our feature set a boolean value indicating whether the output labels of $M_{S}$ and $M_{T}$ are identical or differ. This value, representing the differential testing outcome, is calculated solely based on the two models’ predictions and assigned a value of 1 when both models predict the same class and 0 otherwise. Furthermore, for the reasons presented above, we also feed MetaSel the difference between the logits of two models for each output class, as this may enhance its discriminative power. The difference between $L_{S}$ and $L_{T}$ indicates the magnitude of discrepancies between the prediction of the pre-trained and fine-tuned models for each output class, indicating whether the given input is near the input subspace where the decision boundary has changed.
@@
Notably, we also observed that incorporating representations of target input images derived from a pre-trained VGG-16 [26] model in MetaSel’s feature set did not improve its performance. This may be explained by the fact that information about a given input image is already encoded in the outputs of the logit layer [27]. To validate the contribution of each selected feature to the overall performance of MetaSel, we conduct a comprehensive ablation study and report the results in the subsequent sections.
@@ IV-B Constructing MetaSel’s training set
To train MetaSel, we require a set of inputs from the target distribution ($D_{T}$) whose corresponding MetaSel’s outputs are known. To be able to determine MetaSel’s output, we need to know the following information for each input ($t\in X_{T}$):
@@
To address this, we employed once again ODIN scores [23], to filter out source test inputs that do not align with the target data distribution, $D_{T}$. Source test inputs with ODIN scores lower than a specific threshold are considered OOD and filtered out. To determine an optimized threshold for each $M_{T}$, we rely on the common procedure introduced by Liang et al. [23] that involves using a separate validation set. Following this procedure, we select a threshold that minimizes the False Positive Rate (FPR) at a fixed True Positive Rate (TPR) of 95%, as suggested in the original paper [23]. This threshold ensures that the model retains a high level of accuracy in distinguishing ID inputs from OOD inputs. By applying this filtering mechanism, as illustrated in Figure 2, we ensured that only those source test inputs that satisfy both criteria were included in the MetaSel training set.
@@ V-A Research Questions
RQ4: How does each selected training feature contribute to MetaSel’s performance?
To address this research question, we perform a comprehensive ablation study to evaluate the contribution of each training feature to MetaSel’s effectiveness. Specifically, we construct multiple variants of MetaSel, each excluding one or a group of the original training features. We then compare the performance of the original MetaSel with these variants by analyzing the TRC they achieve. Our goal is to demonstrate that MetaSel, in its original form, maintains superior performance across all variants.
@@ V-B Subjects
We conducted our experiments using a collection of well-known image input sets, including MNIST [28], Fashion-MNIST [28], Cifar-10 [29], and Cifar-100 [29], that have been utilized in numerous empirical studies on DNN testing [30, 3, 31, 19, 32]. MNIST is a set of images of handwritten digits, and Fashion-MNIST (FMNIST) contains images of clothes that are associated with fashion and clothing items. Each one represents a 10-class classification problem and consists of 70,000 grayscale images of a standardized 28x28 size. The training set of each input set includes 60,000 images, while each test set includes 10,000 images. Cifar-10 is another widely used input set that includes a collection of images from 10 different classes (e.g., cats, dogs, airplanes, cars). Furthermore, we selected a more complex input set, Cifar-100, containing 100 classes with 60,000 images total, known for its realistic and feature-rich images. Both Cifar-10 and Cifar-100 are widely used benchmarks in computer vision and contain 32x32 cropped colored images. Each input set is divided into 50,000 images for training and 10,000 images for testing.
@@
In addition to corrupted test sets, we require corrupted versions of our source training sets, i.e., target training set, to perform the fine-tuning process and obtain fine-tuned models. However, Cifar-10-C, Cifar-100-C, MNIST-C, and FMNIST-C datasets only provide corrupted test sets. Therefore, we utilized the original source code and settings employed in the creation of these corrupted test sets to generate the corrupted versions of our source training sets. However, in practical scenarios, the size of the target training set is often limited and significantly smaller than the source training set due to labeling costs. Therefore, we sampled a subset of these corrupted training sets to serve as the target training set, as detailed in the following section. We refer to each fine-tuned DNN model ($M_{T}$) and its corresponding training and test input sets ($Train_{T}$ and $Test_{T}$, respectively) as subjects in our experiments. Consequently, we conducted a comprehensive study across 68 subjects, encompassing four source input sets, seven corruption types, and three levels of corruption severity. To ensure full transparency and reproducibility, all fine-tuned models, along with their corresponding pre-trained counterparts, are publicly available in our replication package.
@@ V-D Evaluation Measure
$\text{TRC (T)}=\frac{Mis_{T}}{min(b,Mis_{total})}$ (5)
where $Mis_{total}$ represents the total number of failures (i.e., mispredicted inputs) in the entire test set, $Mis_{T}$ is the number of failures detected by test subset $T$, and $b$ denotes the selection budget, i.e., size of $T$. Note that the denominator in the TRC formula, as defined in Equation (5), adjusts for budget sizes smaller than $Mis_{total}$. This ensures that TRC accurately evaluates how effective test selection is compared to the ideal performance within the given budget. If the denominator were fixed at $Mis_{total}$, regardless of the budget, achieving a TRC value of 1 would be impossible for any subset selected with $b<Mis_{total}$. For example, with $Mis_{total}$=20 and a selection budget of bb=10, no selected subset could reach a TRC value of 1 due to the inherent budget constraint. The formulation in Equation (5) overcomes this limitation by adapting to the budget, thereby providing a meaningful and general measure of effectiveness across all budgets and regardless of the number of mispredicted inputs.
@@
Figure 3: TRC values for MetaSel and 11 baseline approaches across diverse selection budgets on all subjects with a medium severity level of distribution shift (severity level = 3)
@@
Finally, our primary focus in this context is on evaluating an approach’s capability to identify the maximum number of misclassifications within a specified selection budget. The TRC metric effectively captures this objective. Therefore, we employ TRC as the evaluation metric in our experiments to assess and compare the performance of MetaSel against the baseline methods. We nevertheless report APFD results in our replication package [22] for completeness.
@@ V-E Execution Environment
We conducted our experiments using a cloud computing environment provided by the Digital Research Alliance of Canada [13], utilizing the Cedar cluster with 1352 various Nvidia GPUs (P100 Pascal and V100 Volta) with memory ranging from 12 to 32 GB. Conducting our extensive experiment, which involved deploying MetaSel and all baselines across all subjects and distribution shift severity levels, required approximately 48 days to complete (not accounting for interruptions).
@@ VI-A RQ1: MetaSel’s Performance improvement
[…]
@@
Both DATIS and NNS demonstrate significantly lower TRC medians compared to MetaSel. They also show greater variability in TRC values, introducing the risk of ineffective test selection for some subjects. These findings confirm MetaSel’s superiority and its ability to maintain consistently higher TRC under tight selection budgets, making it the most effective selection approach among all evaluated methods. Moreover, the results highlight the risk of relying on alternative baselines, even the most frequent second-best approaches, like Meta-model, DATIS, or NNS, as they may lead to significantly poor performance. Furthermore, these findings empirically support the core assumption underlying the design of MetaSel—namely, that target test inputs causing behavioral divergence between a fine-tuned model and its pre-trained counterpart are more likely to be misclassified by the fine-tuned model.
@@
Gini Margin Vanilla DSA MDSA LSA TestRank Meta-model DATIS NNS RTS
LeNet5 MNIST 10.45 1.16 0.95 0.74 165.07 1.08 5.40 79.83 9.23 8.98 5.04 141.30
LeNet5 FMNIST 10.64 1.15 0.99 0.97 172.69 1.09 11.03 97.29 11.69 10.11 4.85 154.76
ResNet20 Cifar-10 51.02 2.14 2.02 2.01 167.40 2.98 9.22 92.46 24.81 9.74 5.55 184.64
ResNet152 Cifar-100 277.95 27.14 25.65 25.58 228.40 48.32 63.32 118.07 382.05 91.41 83.25 1282.75
@@ VI-B RQ2: MetaSel’s efficiency
While MetaSel, similar to many baseline approaches, exhibits an increase in execution time as the complexity of the DNN model and the size of the test set grow, it is not the most time-consuming approach. When applied to subjects with relatively simple architectures, such as LeNet-5 or ResNet-20, and input sets with only 10 output classes (i.e., MNIST, FMNIST, or Cifar-10), MetaSel demonstrates efficiency outperforming baselines such as TestRank, DSA, and RTS in terms of execution time. The primary factors driving MetaSel’s execution time are the depth and complexity of the DNN model, as well as the number of output classes. This is primarily due to the requirement for ODIN score calculations for test inputs based on both the pre-trained and fine-tuned models, which requires significantly more computations for forward and backward passes to calculate the ODIN scores on deeper networks such as ResNet152. Additionally, the size of the logit layer, which corresponds to the number of output classes, further increases computational requirements for input sets like Cifar-100. Despite these challenges, MetaSel maintains acceptable efficiency even when applied to deeper networks like ResNet152 and input sets with a larger number of output classes, such as Cifar-100 with 100 classes. Indeed, MetaSel requires approximately 278 seconds to rank the entire test set, a remarkably lower execution time compared to Meta-model, the most viable alternative for practitioners based on effectiveness, which takes around 382 seconds. The increased execution time, combined with the longer initial training phase required by Meta-model, highlights the considerably more time-intensive nature of this method compared to MetaSel. Notably, RTS [20] is the most computationally expensive approach, requiring 4 to 15 times more time than MetaSel to prioritize the entire test set across investigated subjects. This significant overhead primarily stems from its diversity estimation step, which involves computing pairwise similarities between output probability distributions using SSIM for each test input against a sampled subset of training inputs. This computation substantially increases the runtime, particularly for larger input sets. For instance, applying RTS on the Cifar-100 input set takes approximately 1282 seconds.
@@
Answer to RQ2: MetaSel’s efficiency, though lower than many baselines, remains acceptable from a practical standpoint, even for subjects with large input sets featuring numerous output classes, and deep, complex DNN architectures. Its efficiency is even better than some baselines, including the second-best baseline in terms of effectiveness, namely Meta-model. MetaSel is thus a cost-effective approach.